OpenStudy (anonymous):

PLEASE HELP! SOLVE USING: y=mx+b 1. Using the slope-intercept form of a line, find the equation of the line with slope 3/2 and y-intercept –2. Then graph the line on the grid provided and explain how you determined your graph.

OpenStudy (phi):

this question is asking if you can remember what slope-intercept form of a line means they give you the formula y=mx+b which is a big hint. what does the "m" stand for in the formula ?

OpenStudy (anonymous):

mx= 3/2

OpenStudy (phi):

m is the slope for your problem , the slope is 3/2 so m=3/2 just m= 3/2 the x is the "variable" so far you have y= 3/2 x + b

OpenStudy (phi):

in slope-intercept form y= mx+b you have two things (other than the "x") : slope and intercept you know that m is the slope. any guess what b is ?

OpenStudy (anonymous):

-2

OpenStudy (phi):

so using b (the intercept) is -2, your equation is y= (3/2) x - 2

OpenStudy (phi):

to graph it, you need to find two points on the line. plot the two points, and draw a line through the points

OpenStudy (anonymous):

Thank you :)